OpenStudy (anonymous):

The beginning tennis class has twice as many students as the advanced class. The intermediate class has three more students than the advanced class. How many students are in the advanced class if the total enrollment for the three tennis classes is 39? A. 10 B. 15 C. 9 D. 12

OpenStudy (anonymous):

I think its D

OpenStudy (anonymous):

it's C

OpenStudy (anonymous):

2a=b a+3=i b+i+a=39 i suggest checking your math again seeing how i got a different answer, but then again i did this in 2 seconds in my head so i could be wrong

OpenStudy (anonymous):

b = 2a i = a+3 total: 2a + a+3 + a = 4a+3 = 39 a = 9

OpenStudy (texaschic101):

b = beginning class a = advanced class i = intermediate class b + a + 1 = 39 b = 2a i = a + 3 now sub 2a in for b and (a + 3) in for i in the first equation... b + a + i = 39 (2a) + a + (a + 3) = 39 -- combine like terms 4a + 3 = 39 4a = 39 - 3 4a = 36 a = 9 ANSWER : C advanced = 9 intermediate = a + 3 = 9 + 3 = 12 beginning = 2a = 2(9)= 18 check... B + a + i = 39 18 + 9 + 12 = 39 39 = 39 (correct)

OpenStudy (texaschic101):

thank you completeidiot....lol.....luv the name

OpenStudy (anonymous):

Thank You ALL!